- 18 Apr 2011
LBC technical; Mirror temperature modelling: UTC 2011 April 18
Observer(s): A. Rakich, O. Kuhn (LBTO, Tucson)
Telescope Operator: G. Bechetti (LBTO)
Telescope Support: M. Midkiff (LBTO)
Instrument Support: D. Officer (LBTO)
Cycled M1 temperatures while running FPIA continuously. We are attempting to stay on the same circumpolar field all night so as to minimize any elevation changes. Field in Hydra, RA 12 56 44 Dec 66 05 00. The problem with LBC sending incomplete files persisted through the night. Issuetrak 3159. Should have a reasonable set of data here.
M1 temps at start of night: Glass therms SX 6.874; DX 6.718.
All times are UTC
unless otherwise noted.
2:46. Problems with LBC sending incomplete files. See issue trak 3159. Priority increasing to "high".
2:52. FPIA converges in 2 iterations. Doing Rangebal somewhere near our field. Seeing about 2".
3:16 Look at lbcr 031650 for an example of a very bad pupil fit.
4:00. Outside temps and structure aren't that different from the mirror. Now moving to make the mirrors +2 C over ambient, changing setpoints to 2.1.
4:55 just realized that FPIA aborted due to above mentioned issuetrak around 4:37. Restarting FPIA.
5:22 stabalizing mirror temperatures (glass therms) at 7.5 C for SX and 7.2 C for DX (outside air ~ 5.4 C).
5:45 driving the mirror temperatures down now.
6:30 Gino repointing to check wind, pausing our program.
7:05 Back into it, large focus change...
7:55 big jump in seeing (to 3") and associated (and obviously wrong) big jump in Z7 on both SX and DX.
8:42 LBC crashed, power cycled twice, now back in business, another big jump in focus.
9:00 getting mirrors to stabalize at current glass therms (SX 3.4 C, DX 2.8 C).
9:40 now driving back up in temperature (SX 3.112, DX, 2.620).
10:50 now attempting to stabalize at current temperatures. Glass therms: SX, 4.4 C, DX, 3.9 C.
12:00 Crack of dawn, going home.